# Depth-constrained ASV navigation with deep RL and limited sensing

### v1

@@ I INTRODUCTION
In this paper, we address the problem of ASV navigation under depth constraints using RL, where the vehicle must reach a target while avoiding both excessively deep areas and the shoreline, so called unsafe areas. Specifically, we consider the practical case where limited sensing devices are available on the ASV, using a single downward-facing Single Beam Echosounder (SBES) that provides a single depth measurement at the current position. This constraint reflects a growing trend toward low-cost, lightweight, and easily deployable ASVs for environmental monitoring applications, where minimal sensor setups are preferred to reduce complexity and cost—such as in the INTCATCH project, which promotes portable and affordable solutions for water quality assessment  [^1] INTCATCH Project – Development and application of Novel, Integrated Tools for monitoring and managing Catchments. https://www.intcatch.eu. This indirect information requires the agent to infer the locations of unsafe areas without explicit knowledge of their positions. This assumption makes the environment partially observable and significantly limits the agent’s awareness of the surrounding terrain, preventing the agent from inferring a safe direction of motion and thus making the aquatic navigation problem very challenging. Partial observability is commonly addressed through four main approaches: belief state methods [8], which maintain a distribution over possible states; history-based methods [9], such as LSTMs, that use past observations; model-based methods [10], which learn and plan with environment models; and direct observation methods [11], which operate solely on current inputs. Our approach frames partial observability as a belief-state estimation problem, and we employ Gaussian Process (GP) regression to incrementally estimate a spatial belief of the bathymetric map from sparse SBES measurements (Figure 1). While GPs offer principled uncertainty estimates, several studies have shown that predictive variance can saturate in regions with sparse data due to limitations in kernel design and finite-rank approximations [12, holzenkamp2024evaluation, 13]. This leads to overconfident predictions that blur the distinction between observed and inferred areas, compromising the reliability of uncertainty quantification.
@@ II RELATED WORKS
Current research in the domain of aquatic navigation—particularly surface navigation using ASVs—generally falls into three primary categories: Model-based methods, RL-based methods and depth-constrained navigation approaches.
@@ III METHOD
We now introduce the problem of depth-constrained navigation for the ASV, and outline our methodology to solve it, based on RL and GP.
@@ III-A Task Definition
The depth-constrained ASV navigation task (Figure 3) is formulated as an optimization problem, aiming to minimize the distance to a designated target while avoiding unsafe areas, represented by the shoreline and the regions where the water depth exceeds a specified threshold $L_{d}$. The ASV relies on a Global Positioning System (GPS) receiver, an Inertial Measurement Unit (IMU), and a downward-facing SBES, which helps detect the shoreline based on depth variations.
@@ III-B ASV Dynamics
where $X_{\text{tot}}$, $Y_{\text{tot}}$, and $Z_{\text{tot}}$ represent the resultant forces in surge ($x$), sway ($y$), and heave ($z$) directions respectively, while $K_{\text{tot}}$, $M_{\text{tot}}$, and $N_{\text{tot}}$ denote the moments about the roll ($x$), pitch ($y$), and yaw ($z$) axes. The terms $I_{x}$, $I_{y}$, and $I_{z}$ correspond to the principal moments of inertia, with $u$, $v$, $w$ and $p$, $q$, $r$ representing the linear and angular velocities in the body-fixed frame [28].
@@ III-C GP depth estimation
Incorporating only a single depth measurement per time step provides insufficient information about the underlying depth map, making safe navigation challenging—as evidenced by the poor performance of our baseline model w/o GP in the empirical evaluation. To address this limitation, we employ localized GP regression to estimate the depth of nearby locations (Figure 1), leveraging past SBES readings to construct a probabilistic representation of the underwater terrain. This approach enables the vehicle to infer depth variations beyond its immediate measurement, improving decision-making in sparse sensing conditions.
@@
where $\mu_{0}(\mathbf{x})$ is the prior mean function. s For each location $\mathbf{x}^{i}_{t}\in\mathcal{M}(\mathbf{x}_{t})$, the posterior parameters of the Gaussian process are updated by incorporating the new data in a Bayesian manner.
@@ III-E3 Observation space
At each time step $t$, the agent receives a temporal sequence of observations $\{o_{i}\}_{t-T}^{t}$, including key environmental and agent-related variables. Specifically, each observation $o_{i}$ includes the relative target’s position in polar coordinates as $(r_{i},\theta_{i})$, where $r_{i}$ denotes the radial distance and $\theta_{i}$ denotes the angular distance in boat frame; the linear acceleration $\dot{v}_{i}$ and angular acceleration $\dot{\omega}_{i}$; the previous action $a_{i-1}$; the current depth reading $z_{i}$; the estimated depth gradients and the confidence proxy in four primary directions: forward, backward, right and left. Let $\mathbf{x}^{m}_{i}$ be the adjacent position in the main directions to the $\mathbf{x}_{i}$, where $m\in\{f,b,l,r\}$. For each direction, the gradient is computed as:
@@ 4. RQ4
    How does our online GP-based estimation of the local depth map compare to a pre-trained LSTM network [30]?
@@ IV-B Simulation Results
    Privileged model: This model is provided with full access to the true depth map of the lake, serving as an upper bound on performance.
@@
To assess the generalization capabilities (RQ5), we evaluated the best policies for the different methodologies, using 100 additional depth maps that were entirely unseen during training. Each model was evaluated for 1,000 episodes on these unseen maps, with results presented in Table I. The results confirm the evidence for RQ1-4 obtained from the training experiments. Specifically, our model demonstrates the closest performance to the privileged model (86.78% average success rate), increasing the average success rate to 76.26% compared to the model without GP (54.52%). Notably, the LSTM success rate (56.08%) is similar to the case where no depth estimation is available, confirming the crucial role of efficient online GP regression and update, compared to a non-generalizable pre-trained neural model. We also evaluate the efficiency score, i.e., the average success rate over the average episode length, to assess the capability of each policy to solve the task safely and fast. Even in this case, our methodology attains the best result (74.5%), very close to the privileged upper bound (78.1%) and much higher than other competitors. We further evaluate the mean depth break, representing the percentage of episodes where the depth constraint is violated. Our policy attains the second-best result (12.16%), slightly lower than the model without GP (10.98%), though with overlapping variance. This is however balanced by the significantly higher success rate and efficiency of our methodology.
TABLE II: Ablation study results showing the impact of different components.
@@ IV-C Ablation Study
We conducted an ablation study over 1,000 episodes in Table II, to highlight the independent contribution of 3 key items in our methodology: GP estimation, Gradient-Based Extrapolation (explained in III-D) and proxy (Equation 5). Removing the GP module significantly reduced performance, confirming its role in depth-aware navigation. Enabling GP (row 2) improved the baseline, while combining it with Gradient-Based Extrapolation (row 3) further boosted the Success Rate and Efficiency Score. Adding the Variance component (row 4) led to a 2.52% drop in Success Rate and a 1.8% increase in depth violations, likely due to overconfident estimations. Replacing Variance with the Proxy module (row 5) achieved the highest Success Rate (74.14%) and Efficiency Score (0.698) with moderate depth constraints. These results affirm that GP, Gradient-Based Extrapolation, and Proxy are critical for optimal performance.
@@ V CONCLUSION
In this work, we presented a reinforcement learning-based approach for ASV navigation under depth constraints, enabling safe and efficient operation in shallow-water environments with minimal sensor information. Our method integrates Gaussian Process regression to estimate a bathymetric depth map from sparse sonar readings, improving environmental awareness and decision-making. We demonstrated the effectiveness of our approach through sim-to-real transfer, validating its ability to generalize to real-world aquatic conditions. These findings suggest that reinforcement learning, combined with probabilistic modeling, holds significant promises for enhancing autonomous vehicle operation in challenging environments and deserves further exploration. A key limitation of our current approach lies in its reliance on smooth depth estimations from Gaussian Process regression, which may struggle to capture abrupt underwater terrain changes or unmodeled obstacles such as submerged rocks. To address this, future work will explore enhanced sensing strategies—such as incorporating additional sonar sensors—and investigate how the navigation problem can be tackled in a mapless setting, where the ASV must operate without an explicit depth map.

### v2

@@ I INTRODUCTION
In this paper, we address the problem of ASV navigation under depth constraints using RL, where the vehicle must reach a target while avoiding both excessively deep areas and the shoreline, so called unsafe areas. Specifically, we consider the practical case where limited sensing devices are available on the ASV, using a single downward-facing Single Beam Echosounder (SBES) that provides a single depth measurement at the current position. This constraint reflects a growing trend toward low-cost, lightweight, and easily deployable ASVs for environmental monitoring applications, where minimal sensor setups are preferred to reduce complexity and cost, such as in the INTCATCH project, which promotes portable and affordable solutions for water quality assessment  [^1] INTCATCH Project – Development and application of Novel, Integrated Tools for monitoring and managing Catchments. https://www.intcatch.eu. This indirect information requires the agent to infer the locations of unsafe areas without explicit knowledge of their positions. This assumption makes the environment partially observable and significantly limits the agent’s awareness of the surrounding terrain, preventing the agent from inferring a safe direction of motion and thus making the aquatic navigation problem very challenging. Partial observability is commonly addressed through four main approaches: belief state methods [8], which maintain a distribution over possible states; history-based methods [9], such as LSTMs, that use past observations; model-based methods [10], which learn and plan with environment models; and direct observation methods [11], which operate solely on current inputs. Our approach frames partial observability as a belief-state estimation problem, and we employ Gaussian Process (GP) regression to incrementally estimate a spatial belief of the bathymetric map from sparse SBES measurements (Figure 1). While GPs offer principled uncertainty estimates, several studies have shown that predictive variance can saturate in regions with sparse data due to limitations in kernel design and finite-rank approximations [12, 13, 14]. This leads to overconfident predictions that blur the distinction between observed and inferred areas, compromising the reliability of uncertainty quantification.
@@ II RELATED WORKS
Current research in the domain of aquatic navigation, particularly surface navigation using ASVs, generally falls into three primary categories: Model-based methods, RL-based methods and depth-constrained navigation approaches.
@@ III METHOD
We now introduce the depth-constrained ASV navigation problem and outline our RL- and GP-based solution.
@@ III-A Task Definition
The depth-constrained ASV navigation task (Figure 3) is formulated as an optimization problem, aiming to minimize the distance to a designated target while avoiding unsafe areas, represented by the shoreline and the regions where the water depth exceeds a specified threshold $L_{d}$. The ASV relies only on a downward-facing SBES, which helps detect the shoreline based on depth variations.
@@ III-B ASV Dynamics
where $X_{\text{tot}}$, $Y_{\text{tot}}$, and $Z_{\text{tot}}$ represent the resultant forces in surge ($x$), sway ($y$), and heave ($z$) directions respectively, while $K_{\text{tot}}$, $M_{\text{tot}}$, and $N_{\text{tot}}$ denote the moments about the roll ($x$), pitch ($y$), and yaw ($z$) axes. The terms $I_{x}$, $I_{y}$, and $I_{z}$ correspond to the principal moments of inertia, with $u$, $v$, $w$ and $p$, $q$, $r$ representing the linear and angular velocities in the body-fixed frame.
@@ III-C GP depth estimation
Incorporating only a single depth measurement per time step provides insufficient information about the underlying depth map, making safe navigation challenging, as evidenced by the poor performance of our baseline model w/o GP in the empirical evaluation. To address this limitation, we employ localized GP regression to estimate the depth of nearby locations (Figure 1), leveraging past SBES readings to construct a probabilistic representation of the underwater terrain. This approach enables the vehicle to infer depth variations beyond its immediate measurement, improving decision-making in sparse sensing conditions.
@@
where $\mu_{0}(\mathbf{x})$ is the prior mean function. For each location $\mathbf{x}^{i}_{t}\in\mathcal{M}(\mathbf{x}_{t})$, the posterior parameters of the Gaussian process are updated by incorporating the new data in a Bayesian manner.
@@ III-E3 Observation space
At each time step $t$, the agent receives a temporal sequence of observations $\{o_{i}\}_{t-T}^{t}$, including key environmental and agent-related variables. Specifically, each observation $o_{i}$ includes the relative target’s position in polar coordinates as $(r_{i},\theta_{i})$, where $r_{i}$ denotes the radial distance and $\theta_{i}$ denotes the angular distance in boat frame; the linear acceleration $\dot{u}_{i}$ and angular acceleration $\dot{\omega}_{i}$; the previous action $a_{i-1}$; the current depth reading $z_{i}$; the estimated depth gradients and the confidence proxy in four primary directions: forward, backward, right and left. Let $\mathbf{x}^{m}_{i}$ be the adjacent position in the main directions to the $\mathbf{x}_{i}$, where $m\in\{f,b,l,r\}$. For each direction, the gradient is computed as:
@@ 4. RQ4
    How does our online GP-based estimation of the local depth map compare to a pre-trained LSTM network?
@@ IV-B Simulation Results
    Privileged model: This model has privileged access to true depth gradients, bypassing estimation and serving as a performance upper bound.
@@
To evaluate generalization (RQ5), we tested the best policies on 100 unseen depth maps, running 1,000 episodes per model (see Table I). Results align with findings from RQ1–4. Our model performs closest to the privileged one (76.26% vs. 86.78% success), significantly outperforming the w/o GP baseline (54.52%) and LSTM (56.08%), highlighting the importance of online GP regression over static neural models. We also assessed the efficiency score, where our method again ranks highest (0.745), close to the privileged upper bound (0.781). For mean depth break (percentage of episodes violating depth constraints), our approach ranks second (12.16%), slightly behind the no-GP model (10.98%) but with significantly higher success and efficiency.
TABLE II: Ablation study results showing the impact of different components. The last line represents our full model. Row 2 uses techniques proposed by [15], while row 4 combines techniques from both [15] and [16].
@@ IV-C Ablation Study
We conducted an ablation study over 1,000 episodes in Table II, to highlight the independent contribution of 3 key items in our methodology: GP estimation, Gradient-Based Extrapolation (explained in III-D) and proxy (Equation 5). Removing the GP module significantly reduced performance, confirming its role in depth-aware navigation. Enabling GP (row 2) [15] improved the baseline, while combining it with Gradient-Based Extrapolation [16] (row 3) further boosted the Success Rate and Efficiency Score. Adding the Variance component (row 4) led to a 2.52% drop in Success Rate and a 1.8% increase in depth violations, likely due to overconfident estimations. Replacing Variance with the Proxy module (row 5) achieved the highest Success Rate (74.14%) and Efficiency Score (0.698) with moderate depth constraints. These results affirm that GP, Gradient-Based Extrapolation, and Proxy are critical for optimal performance.
IV-D Real-World ASV Navigation
To evaluate the transferability of our policy from simulation to the real world, we deployed it on a physical ASV built using off-the-shelf components, including a 3D printed frame, two underwater thrusters, one SBES sensor, and a Raspberry Pi 3. The ASV was tasked with reaching randomly generated target points in a real aquatic environment. We conducted five experiments, with the ASV successfully reaching the target in four out of five runs. The only failure was due to a depth constraint violation, caused by fluctuations in the depth sensor’s sampling frequency. The efficiency score (ES) was 0.530; vertical and heading stability metrics (VS and HS) were not computed due to the lack of an onboard IMU. While the policy showed generally reliable performance, we observed fluctuations in control frequency stemming from the Raspberry Pi’s limited processing power, occasionally leading to degraded control. Communication delays, which were not modeled during training, also contributed to the sim-to-real performance gap.
@@ V CONCLUSION
In this work, we presented a reinforcement learning-based approach for ASV navigation under depth constraints, enabling safe and efficient operation in shallow-water environments with minimal sensor information. Our method integrates Gaussian Process regression to estimate a bathymetric depth map from sparse sonar readings, improving environmental awareness and decision-making.
We validated our approach through extensive simulation experiments and demonstrated its transferability to real-world aquatic scenarios. The policy achieved reliable real-world performance despite limited onboard computation and unmodeled environmental factors, highlighting its robustness and practical potential.
A key limitation lies in the reliance on smooth depth estimations from GP regression, which may struggle with abrupt terrain changes or unmodeled obstacles such as submerged rocks. To address this, future work will also explore enhanced sensing strategies (e.g., additional sonar units) and investigate mapless navigation approaches, where the ASV operates without an explicit depth map. To improve real-world performance, especially under limited hardware and communication constraints, future work will explore domain randomization and adaptation techniques to enhance robustness and reduce the sim-to-real gap.